• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Solvignon

    Log entry time 00:01:51 on April13,2007

    Entry number 198814

    keyword=Shift summary

    16:00 Start shift on 3He (Q2 = 55 fm-2)

    16:20 The target vacuum is increasing so we lowered the current to
    50uA and now it looks stable. Now we are ramping up the current again and watch the fluctuation in the target vacuum.

    20:08 changed raster size from 5x2 (MCC) to 5x3 (MCC)


    00:00 Shift finished with run 2346 underway.

    ------------
    run# 2369: 15:57-->16:13, 100uA, 1.4M
    run# 2370: 16:15-->16:38, 50-100uA, 1.4M (stabilizing vacuum of the target)
    run# 2371: 16:40-->17:05, 90-100uA, 1.4M
    run# 2372: 17:18-->17:38, 95uA, 1.4M (instable beam)
    run# 2373: 17:38-->17:56, 95uA, 1.4M
    run# 2374: 17:57-->18:16, 95uA, 1.4M (ps6=4)
    run# 2375: 18:17-->18:26, 95uA, 0.6M (ps6=5, ps1=1000)
    run# 2376: 18:27-->18:46, 95uA, 1.4M (ps6=20)
    run# 2377: 18:47-->19:05, 95uA, 1.4M
    run# 2378: 19:05-->19:23, 95uA, 1.4M
    run# 2379: 19:24-->19:44, 95uA, 1.4M
    run# 2380: 19:44-->20:06, 95uA, 1.4M
    run# 2381: 20:07-->20:28, 95uA, 1.4M
    run# 2382: 20:28-->20:31, 95uA, 0.2M
    run# 2383: 20:33-->20:53, 95uA, 1.4M Changed raster size to 8x3 (MCC)
    run# 2384: 20:54-->21:26, 95uA, 1.4M (ps3=2)
    run# 2385: 21:28-->21:58, 95uA, 1.4M
    run# 2386: 21:59-->22:31, 95uA, 1.4M
    run# 2387: 22:31-->23:03, 95uA, 1.4M
    run# 2388: 22:03-->23:46, 95uA, 1.4M
    run# 2389: 23:46-->